OpenStudy (anonymous):

Add or Subtract. Please explain how to do the process so I can understand it!! 7a^2 - 10a^2 + 9a

OpenStudy (accessdenied):

Do you understand the process of distributing? Like, a(b + c) = ab + ac? What I would like to propose for understanding this problem is factoring, which is the reverse process of distributing. (ab + ac = a(b+c))

OpenStudy (anonymous):

I sort of understand it, it's been a while since I've done it.

OpenStudy (accessdenied):

Would it make sense if I said that a^2* (7-10) = 7a^2 - 10a^2 then?

OpenStudy (anonymous):

No not really, sorry lol, i'm really not that great at polynomials and monomials...

OpenStudy (accessdenied):

If we can agree that these are the same thing, then we can say that 7a^2 - 10a^2 = a^2 * (7-10), or (7 - 10)a^2 7a^2 - 10a^2 + 9a = (7 - 10)a^2 + 9a

OpenStudy (anonymous):

oh! I get that now.

OpenStudy (accessdenied):

 I think that's the best way to illustrate how adding/subtracting like-terms works out. If they have the same variable and exponent, then it may be factored out and all that remains are the #'s in front of them. If we tried to add 10a^2 + 9a, factoring = (10a + 9)a, we can't pull out another a to add anything so it fails. :)